OpenStudy (anonymous):

find the area of the triangle with vertices at (2,3,5) ,Q(4,2,1),R(3,6,4)

OpenStudy (anonymous):

This one is quite long. I might have to do it on paper.

OpenStudy (anonymous):

use draw

OpenStudy (anonymous):

P(2, 3, 5) Q(4, 2, 1) R(3, 6, 1) \[a = PQ = \sqrt{(4-2)^{2} + (2 - 3)^{2} + (1 - 5)^{2}} \] \[= \sqrt{2^{2} + (-1)^{2} + (-4)^{2}}\] \[= \sqrt{21}\] \[b = QR = \sqrt{(3-4)^{2} + (6-2)^{2} + (1 - 1)^{2}}\] \[\sqrt{(-1)^{2} + 4^{2}+ 0}\] \[\sqrt{?}\] \[c = RP = \sqrt{(2-3)^{2} + (6-3)^{2} + (1-5)^{2}}\] \[= \sqrt{(-1)^{2} + 3^{2} + (-4)^{?}}\] \[= \sqrt{26}\] Now you just need to convert from surds using Heron's formula and it's a cakewalk from there.

OpenStudy (anonymous):

\[b = QR = \sqrt{17}\]

OpenStudy (anonymous):

Let me know if you get stuck with the rest of the question and I'll walk you through it.

OpenStudy (anonymous):

but remember that area=1/2bh

OpenStudy (anonymous):

am talking about area of the triangle not vector

OpenStudy (anonymous):

There are different ways you can do this. My method is to find the length of each side, convert them to decimal, then use Heron's formula to find the area.

OpenStudy (anonymous):

You could also take the cross product of two of the points, then define the origin and work from that.

OpenStudy (anonymous):

please use the second method and let me see. Thanks

OpenStudy (anonymous):

go further by using the heron formular

OpenStudy (anonymous):

Firstly: \[a = \sqrt{21} = 4.583\] \[b = \sqrt{17} = 4.123\] \[c = \sqrt{26} = 5.099\] Heron's formula is denoted as \[A = \sqrt{(s(s-a)(s-b)(s-c))}\] where \[s = (a+b+c)/2\] Calculate s using the above equation, then use the area equation to find the area of the triangle.

OpenStudy (anonymous):

http://mathworld.wolfram.com/TriangleArea.html

OpenStudy (anonymous):

I am not sure if this form of Heron's formula works in 3 dimensions, since the 3D version of Pythagorean theprem is different http://www.jstor.org/discover/10.2307/3619577?uid=3738256&uid=2129&uid=2&uid=70&uid=4&sid=47698963103007

OpenStudy (anonymous):

That's for a pyramid with 4 vertices, this is for a triangle lying inside a 3D plane.

OpenStudy (anonymous):

i got A has 80.30

OpenStudy (anonymous):

what is the formular for area

OpenStudy (anonymous):

 can i do it this way